# The metaplectic semigroup and its applications to time-frequency analysis and evolution operators

### v1

@@ Abstract.
We develop a systematic analysis of the metaplectic semigroup $\operatorname*{Mp}_{+}(d,\mathbb{C})$ associated with positive complex symplectic matrices, a notion introduced almost simultaneously and independently by Hörmander, Brunet, Kramer, and Howe, thereby extending the classical metaplectic theory beyond the unitary setting. While the existing literature has largely focused on propagators of quadratic evolution equations, for which results are typically obtained via Mehler formulas, our approach is operator-theoretic and symplectic in spirit and adapts techniques from the standard metaplectic group $\operatorname*{Mp}(d,\mathbb{R})$ to a substantially broader framework that is not driven by differential problems or particular propagators. This point of view provides deeper insight into the structure of the metaplectic semigroup, and allows us to investigate its generators, polar decomposition, and intertwining relations with complex conjugation and with the Wigner distribution. We then exploit these structural results to characterize, from a metaplectic perspective, classes of time-frequency representations satisfying prescribed structural properties. Finally, we discuss further implications for parabolic equations with complex quadratic Hamiltonians, we study the boundedness of their propagators on modulation spaces, we obtain estimates in time of their operator norms. Finally, we apply our theory to the study of propagation of Wigner singularities.
@@ 1. Introduction
Motivated by applications to time-frequency analysis and to evolution equations, the aim of this work is to present the metaplectic semigroup $\operatorname*{Mp}_{+}(d,\mathbb{C})$ in the language of modern harmonic analysis, and to develop a theory analogous to that of the classical metaplectic group $\operatorname*{Mp}(d,\mathbb{R})$. In contrast to most of the existing literature, where it is presented as a semigroup of bounded operators on the Fock space, we study the metaplectic semigroup as a semigroup of operators bounded on $L^{2}(\mathbb{R}^{d})$, in the tradition of Hörmander. In the first three subsections of this Introduction, we review known results about the basics of $\operatorname*{Mp}(d,\mathbb{R})$. Applications to time-frequency analysis and equations of Schrödinger type are recalled. We emphasize that the extension to the “complex” case appears as a natural project for both the issues. Finally, Subsection 1.4 of this Introduction presents our setting of $\operatorname*{Mp}_{+}(d,\mathbb{C})$, with the main results providing a unitary treatment of time-frequency and PDEs problems.
@@ 1.1. The metaplectic representation
where we may assume for simplicity $u_{0}\in\mathcal{S}(\mathbb{R}^{d})$, and $a$ is the Weyl quantization of a real quadratic form $a(z)=\mathcal{Q}z\cdot z$, $z\in{\mathbb{R}^{2d}}$ and $\mathcal{Q}=\mathcal{Q}^{\top}$, that is
@@
If $f=g$, we write $Wf=W(f,f)$, the Wigner distribution of $f$. Then, if $U_{t}\in\operatorname*{Sp}(d,\mathbb{R})$ describes the Hamiltonian flow of (1), the aforementioned cover of the symplectic group maps $e^{it\mathrm{Op}^{\mathrm{w}}(a)}\in\operatorname*{Mp}(d,\mathbb{R})\mapsto U_{t}\in\operatorname*{Sp}(d,\mathbb{R})$ for every $t\in\mathbb{R}$, and these evolution operators read, for every $t\in\mathbb{R}$, as linear symplectic transformations of the phase-space:
@@
is not only satisfied for every $f,g\in L^{2}(\mathbb{R}^{d})$, but it is also an alternative definition of metaplectic operators and of their projection onto $\operatorname*{Sp}(d,\mathbb{R})$ [56]. Indeed, in a more algebraic fashion, the metaplectic operators associated to a symplectic matrix $U\in\operatorname*{Sp}(d,\mathbb{R})$ are classically defined, up to a phase factor, through the intertwining relation with the Schrödinger representation of the Heisenberg group
@@
we refer to Section 2 below for the details. For the purpose of this discussion, the phase factor $e^{2\pi i\tau}$ plays a minor role and will therefore write $\rho(z)$ for the rest of the Introduction, when it is convenient. This dualism between operators and matrices is crucial: the aforementioned cover map is indeed a homomorphism, and this considerably simplifies the analysis of metaplectic operators by reducing operator calculus to linear algebra. This allows properties of metaplectic operators to be inferred and interpreted directly from the structure of their projection. Conversely, the mapping associating to each $U\in\operatorname*{Sp}(d,\mathbb{R})$ the corresponding metaplectic operator, defined up to a phase factor, is a projective representation known as the metaplectic representation.
@@
In (8) and (9) the correct phase factors are omitted for simplicity of the discussion. We leave the details to Section 2. The interest in the metaplectic representation has increased in the last decade, as it is witnessed by the recent contributions of several authors, see e.g. [4, 24, 28, 33, 48].
@@
is the partial Fourier transform with respect to the frequency variables, and $\mathfrak{T}_{E_{w}}F(x,y)=F(x+y/2,x-y/2)$. The identity (10) also extends uniquely the Wigner distributions to $f,g\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$. Both these operators are in $\operatorname*{Mp}(2d,\mathbb{R})$, and so is their composition. The Wigner distribution shares this feature with many other popular time-frequency representations, such as the short-time Fourier transform (STFT)
@@
with the result that time-frequency analysis is strictly woven with metaplectic operators’ theory. Similarly to the Wigner distribution, the STFT can be generalized to $f,g\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$.
@@
By working in this setting, we can exploit the deep relation of the metaplectic group with symplectic matrices to express and explain the properties of an $\mathcal{U}$-Wigner distribution in terms of the structure of the projection $\mathcal{U}$. Nowadays, metaplectic Wigner distributions are a well-established subject of study in the time-frequency analysis community, their applications spanning from pseudodifferential calculus [3, 23] to frame theory [13, 21], uncertainty principles [34], quantum physics [22] and signal processing [33, 46, 51]. In [35], it has been proven that, under minimal continuity assumptions, a time-frequency representation $Q:\mathcal{S}(\mathbb{R}^{d})\times\mathcal{S}(\mathbb{R}^{d})\to\mathcal{S}^{\prime}({\mathbb{R}^{2d}})$ satisfies
@@
A time-frequency representation $Q:L^{2}(\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d})\to L^{2}({\mathbb{R}^{2d}})$ is covariant if
(16) $Q(\rho(z)f,\rho(z)g)=Q(f,g)(\cdot-z).$
@@ 1.3. The metaplectic semigroup $\operatorname*{Mp}_{+}(d,\mathbb{C})$: former contributions
The four leading figures who investigated this complex extension of the metaplectic group are Brunet and Kramer [9, 8], Howe [43], and Hörmander [40, 41]. Although they worked independently, they all arrived at the same mathematical object. Brunet and Kramer coined the term metaplectic semigroup, which was also used later by Hörmander, while Howe defined the oscillator semigroup. Howe, Brunet and Kramer argued at the level of the Fock space, where the existence of a reproducing kernel facilitated the definition of operators in terms of their Schwartz kernels (as operators on the Fock space). These two approaches were shown to be equivalent by Hilgert in 1989 [37]. On the contrary, Hörmander’s semigroup of positive symplectic linear maps consists essentially of operators with Gaussian kernel, which form a double cover of a semigroup of positive complex symplectic matrices, see Definition 3.1 in Section 3.1. The equivalence between Hörmander’s theory and Brunet-Kramer-Howe’s theory is discussed for the benefit of the reader in Remark 3.6 below. It can be either inferred by comparing the positivity conditions [41, Equation (5.10)] and [8, Page 214], or by comparing [41, Proposition 5.10] with [9, Theorem 2.1] and [40, Section 4] with [9, Section V]. We devote Section 3 to reviewing the definition and structure of the metaplectic semigroup, denoted here by $\operatorname*{Mp}_{+}(d,\mathbb{C})$, and of its projection onto the subsemigroup of positive symplectic matrices $\operatorname*{Sp}_{+}(d,\mathbb{C})$.
@@
with $\mathcal{Q}^{\top}=\mathcal{Q}$ is a complex quadratic form with positive semi-definite real part. The solution $u(t,x)=e^{-it\mathrm{Op}^{\mathrm{w}}(a)}u_{0}(x)$ is expressed by the generalized Mehler formula proven by Hörmander [41, Theorem 4.1].
The analysis of these problems has initiated a long series of recent and well-celebrated works, extending far beyond Hörmander’s original contributions. Building on Hörmander’s analysis of quadratic operators, Hitrik and Pravda-Starov extended his framework to the non-elliptic setting in [39] and further derive smoothing properties of heat semigroups, spectral discreteness under suitable ellipticity on this space, and the large-time behavior of the generated contraction semigroups. Importantly, they discover the singular space associated to a complex quadratic form, which describes geometrically the (symplectic orthogonal complement of the) directions where the corresponding heat semigroup is smoothing. Microlocal analysis of problems of the form (1) with complex quadratic Hamiltonians was subsequently developed in a series of works [53, 52] and, more recently, in [50, 60], within the framework of propagation of singularities, aimed at encoding the microlocal singularities of solutions of evolution equations of the form (35) below and relate them with those of the initial datum, inside the singular space.
In [2] the authors derive a polar decomposition for one-parameter semigroups describing the propagators of the corresponding problem (1). Particularly noteworthy are the recent contributions of Hitrik, Pravda-Starov, and Viola [38], Viola [58, 59], and Alphonse and Bernier [1, 5], which study the smoothing and subelliptic properties of quadratic non-selfadjoint differential operators, once again relying on the geometric structure of the singular space. A more time-frequency analysis approach is instead developed by Trapasso in [57], where he provides off-diagonal decay estimates for the Gabor matrices of these one-parameter semigroups. It seems natural to develop an operator-theoretic treatment of the general propagator $e^{-it\mathrm{Op}^{\mathrm{w}}(a)}$, with $a(z)$ satisfying (17), in the lines of a metaplectic calculus. This allows applications to the so-called Wigner wave front set and to the boundedness in time-frequency spaces.
@@ 1.4.1. Generators
varying $E\in\operatorname*{GL}(d,\mathbb{R})$ and $Q$ real and symmetric. By allowing $E$ and $Q$ to be complex, we obtain the generators of $\operatorname*{Sp}(d,\mathbb{C})$. However, constructing a projective representation associating to every $S\in\operatorname*{Sp}(d,\mathbb{C})$ a group of operators bounded on $L^{2}(\mathbb{R}^{d})$, which agrees with the metaplectic representation on $\operatorname*{Sp}(d,\mathbb{R})$, is impossible. This reflects the fact that, unlike the real case, problems (1) with complex quadratic Hamiltonians are generally defined only for $t\geq 0$. Indeed, $\operatorname*{Sp}_{+}(d,\mathbb{C})$ forms a semigroup rather than a group, and the question of its generators arises naturally.
@@ 1.4.2. Polar decomposition and its applications
The polar decomposition derived in [2] for the propagators of problems (1) with complex quadratic Hamiltonians is actually a very general result holding for every metaplectic operator, see Lemma 4.4. This yields the algebraic characterization of $\operatorname*{Mp}_{+}(d,\mathbb{C})$ as isomorphic to $\operatorname*{Mp}(d,\mathbb{R})\times\operatorname*{Mp}_{0}(d,\mathbb{C})$, where $\operatorname*{Mp}_{0}(d,\mathbb{C})$ is the subset of metaplectic operators that are positive and self-adjoint, and to a further characterization of the operators in $\operatorname*{Mp}_{0}(d,\mathbb{C})$ as the $\operatorname*{Mp}(d,\mathbb{R})$-conjugates of an explicit subsemigroup of metaplectic operators, here called atomic metaplectic contractions, see Section 3.2 and Corollary 4.5. Let us briefly summarize these results.
@@
can hold for $M\in\operatorname*{GL}(2d,\mathbb{R})$ if and only if the corresponding operator $T$ is either a metaplectic operator or the composition of a metaplectic operator with complex conjugation, as it was proven by Dias and Prata [25].
@@
Since the polar decomposition of $\widehat{S}$ is unique, formula (23) provides an unambiguous separation between the contribution of what can be regarded as a sort of real part of $\widehat{S}$, namely $\widehat{U}\in\operatorname*{Mp}(d,\mathbb{R})$, and that of its imaginary part, represented by $\widehat{Z}\in\operatorname*{Mp}_{0}(d,\mathbb{C})$. However inappropriate, these phraseologies provide a clear picture of how the positive self-adjoint operator $\widehat{A}$, absent for metaplectic operators with real projections, affects the phase-space concentration of $\widehat{S}$. Stated differently, when $\widehat{U}\in\operatorname*{Mp}(d,\mathbb{R})$ we recover the classical symplectic covariance property of the Wigner distribution (5), whereas in the general case (23) yields an extension in which the action of $\widehat{U}$ is disentangled from the additional contribution induced by $\widehat{Z}$.

### v2

@@ Abstract.
We develop a systematic analysis of the metaplectic semigroup $\operatorname*{Mp}_{+}(d,\mathbb{C})$ associated with positive complex symplectic matrices, a notion introduced almost simultaneously and independently by Hörmander, Brunet, Kramer, and Howe, thereby extending the classical metaplectic theory beyond the unitary setting. While the existing literature has largely focused on propagators of quadratic evolution equations, where integral representations are available in specific cases thanks to Mehler formulas, our approach is operator-theoretic and symplectic in spirit and adapts techniques from the standard metaplectic group $\operatorname*{Mp}(d,\mathbb{R})$ to a substantially broader framework that is not driven by differential problems or particular propagators. This point of view provides deeper insight into the structure of the metaplectic semigroup, and allows us to investigate its generators, polar decomposition, and intertwining relations with complex conjugation and with the Wigner distribution. We then exploit these structural results to characterize, from a metaplectic perspective, classes of time-frequency representations satisfying prescribed structural properties. Finally, we discuss further implications for evolution equations with complex quadratic Hamiltonians, we study the boundedness of their propagators on modulation spaces, we obtain estimates in time of their operator norms. Finally, we apply our theory to the study of the propagation of concentration of Wigner distributions.
@@ 1. Introduction
Motivated by applications to time-frequency analysis and to evolution equations, the aim of this work is to present the metaplectic semigroup $\operatorname*{Mp}_{+}(d,\mathbb{C})$ in the language of modern harmonic analysis, and to develop a theory analogous to that of the classical metaplectic group $\operatorname*{Mp}(d,\mathbb{R})$. In contrast to most of the existing literature, where it is presented as a semigroup of bounded operators on the Fock space, we study the metaplectic semigroup as a semigroup of operators bounded on $L^{2}(\mathbb{R}^{d})$, in the spirit of Hörmander. In the first three subsections of this Introduction, we review known results about the basics of $\operatorname*{Mp}(d,\mathbb{R})$. Applications to time-frequency analysis and equations of Schrödinger type are recalled. We emphasize that the extension to the “complex” case appears as a natural project for both the issues. Finally, Subsection 1.4 of this Introduction presents our setting of $\operatorname*{Mp}_{+}(d,\mathbb{C})$, with the main results providing a unitary treatment of time-frequency and PDEs problems.
@@ 1.1. The metaplectic representation
where $u_{0}\in L^{2}(\mathbb{R}^{d})$, and $a$ is the Weyl quantization of a real quadratic form $a(z)=\mathcal{Q}z\cdot z$, $z\in{\mathbb{R}^{2d}}$ and $\mathcal{Q}=\mathcal{Q}^{\top}$, that is
@@
If $f=g$, we write $Wf=W(f,f)$, the Wigner distribution of $f$. Then, if $U_{t}\in\operatorname*{Sp}(d,\mathbb{R})$ describes the canononical transformation associated with the Hamiltonian $a$ in (1), the aforementioned cover of the symplectic group maps $e^{it\mathrm{Op}^{\mathrm{w}}(a)}\in\operatorname*{Mp}(d,\mathbb{R})\mapsto U_{t}\in\operatorname*{Sp}(d,\mathbb{R})$ for every $t\in\mathbb{R}$, and these evolution operators read, for every $t\in\mathbb{R}$, as linear symplectic transformations of the phase-space:
@@
is not only satisfied for every $f,g\in L^{2}(\mathbb{R}^{d})$, but it also characterizes metaplectic operators and of their projection onto $\operatorname*{Sp}(d,\mathbb{R})$ [54]. In a more algebraic fashion, the metaplectic operators associated to a symplectic matrix $U\in\operatorname*{Sp}(d,\mathbb{R})$ are classically defined, up to a phase factor, through the intertwining relation with the Schrödinger representation of the Heisenberg group
@@
we refer to Section 2 below for the details. For the purpose of this discussion, the phase factor $e^{2\pi i\tau}$ plays a minor role and we will therefore write $\rho(z)$ for the rest of the Introduction, when it is convenient. This dualism between operators and matrices is crucial: the aforementioned cover map is indeed a homomorphism, and this considerably simplifies the analysis of metaplectic operators by reducing operator calculus to linear algebra. This allows properties of metaplectic operators to be inferred and interpreted directly from the structure of their projection. Conversely, the mapping associating to each $U\in\operatorname*{Sp}(d,\mathbb{R})$ the corresponding metaplectic operator, defined up to a phase factor, is a projective representation known as the metaplectic representation.
@@
In (8) the correct phase factors are omitted for simplicity of the discussion. We leave the details to Section 2. The interest in the metaplectic representation has increased in the last decade, as it is witnessed by the recent contributions of several authors, see e.g. [4, 24, 28, 31, 46].
@@
is the partial Fourier transform with respect to the frequency variables, and $\mathfrak{T}_{E_{w}}F(x,y)=F(x+y/2,x-y/2)$. Both the operators $\mathcal{F}_{2}$ and $\mathfrak{T}_{E_{w}}$ are in $\operatorname*{Mp}(2d,\mathbb{R})$, and so is their composition. The Wigner distribution shares this feature with many other popular time-frequency representations, such as the short-time Fourier transform (STFT)
@@
with the result that time-frequency analysis is strictly woven with metaplectic operators’ theory. Note that, in this fashion, the definitions of the Wigner distribution and of the STFT can also be generalized to $f,g\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$.
@@
By working in this setting, we can exploit the deep relation of the metaplectic group with symplectic matrices to express and explain the properties of an $\mathcal{U}$-Wigner distribution in terms of the structure of the projection $\mathcal{U}$. Nowadays, metaplectic Wigner distributions are a well-established subject of study in the time-frequency analysis community, their applications spanning from pseudodifferential calculus [3, 23] to frame theory [12, 20], uncertainty principles [32], quantum physics [21] and signal processing [31, 44, 49]. In a recent unpublished work  [^1] See K. Gröchenig and I. Shafkulovska. More uncertainty principles for metaplectic time-Frequency representations. arXiv:2503.13324, 2025., it has been proven that, under minimal continuity assumptions, a time-frequency representation $Q:\mathcal{S}(\mathbb{R}^{d})\times\mathcal{S}(\mathbb{R}^{d})\to\mathcal{S}^{\prime}({\mathbb{R}^{2d}})$ satisfies
@@
A quadratic time-frequency representation $Q:L^{2}(\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d})\to L^{2}({\mathbb{R}^{2d}})$ is covariant if
(16) $Q(\pi(z)f,\pi(z)g)=Q(f,g)(\cdot-z),$
where $\pi(x,\xi)f(y)=e^{2\pi i\xi\cdot y}f(y-x)$.
@@ 1.3. The metaplectic semigroup $\operatorname*{Mp}_{+}(d,\mathbb{C})$: former contributions
The four leading figures who investigated this complex extension of the metaplectic group are Brunet and Kramer [9, 8], Howe [41], and Hörmander [38, 39]. Although they worked independently, they all arrived at the same mathematical object. Brunet and Kramer coined the term metaplectic semigroup, which was also used later by Hörmander, while Howe defined the oscillator semigroup. Howe, Brunet and Kramer argued at the level of the Fock space, where the existence of a reproducing kernel facilitated the definition of operators in terms of their Schwartz kernels (as operators on the Fock space). These two approaches were shown to be equivalent by Hilgert in 1989 [35]. On the contrary, Hörmander’s semigroup of positive symplectic linear maps consists essentially of operators on $L^{2}(\mathbb{R}^{d})$ with Gaussian kernel, which form a double cover of a semigroup of positive complex symplectic matrices, see Definition 3.1 in Section 3.1. The equivalence between Hörmander’s theory and Brunet-Kramer-Howe’s theory is discussed for the benefit of the reader in Remark 3.6 below. It can be either inferred by comparing the positivity conditions [39, Equation (5.10)] and [8, Page 214], or by comparing [39, Proposition 5.10] with [9, Theorem 2.1] and [38, Section 4] with [9, Section V]. We devote Section 3 to reviewing the definition and structure of the metaplectic semigroup, denoted here by $\operatorname*{Mp}_{+}(d,\mathbb{C})$, and of its projection onto the subsemigroup of positive symplectic matrices $\operatorname*{Sp}_{+}(d,\mathbb{C})$.
@@
with $\mathcal{Q}^{\top}=\mathcal{Q}$ a complex quadratic form with positive semi-definite imaginary part. Outside of the caustics, the solution $u(t,x)=e^{-it\mathrm{Op}^{\mathrm{w}}(a)}u_{0}(x)$ can be expressed by Mehler formulas, proven by Hörmander [39, Theorem 4.1].
The analysis of these problems has initiated a long series of recent and well-celebrated works, extending far beyond Hörmander’s original contributions. Building on Hörmander’s analysis of quadratic operators, Hitrik and Pravda-Starov extended his framework to the non-elliptic setting in [37] and further derive smoothing properties of heat semigroups, spectral discreteness under suitable ellipticity on this space, and the large-time behavior of the generated contraction semigroups. Importantly, they discover the singular space associated to a complex quadratic form, which describes geometrically the (symplectic orthogonal complement of the) directions where the corresponding heat semigroup is smoothing. Microlocal analysis of problems of the form (1) with complex quadratic Hamiltonians was subsequently developed in a series of works [51, 50] and, more recently, in [48, 58], within the framework of propagation of singularities, aimed at encoding the microlocal singularities of solutions of quadratic evolution equations and relate them with those of the initial datum, inside the singular space.
In [2] the authors derive a polar decomposition for one-parameter semigroups describing the propagators of the corresponding problem (1). Particularly noteworthy are the recent contributions of Hitrik, Pravda-Starov, and Viola [36], Viola [56, 57], and Alphonse and Bernier [1, 5], which study the smoothing and subelliptic properties of quadratic non-selfadjoint differential operators, once again relying on the geometric structure of the singular space. A more time-frequency analysis approach is instead developed by Trapasso in [55], where he provides off-diagonal decay estimates for the Gabor matrices of these one-parameter semigroups. It seems natural to develop an operator-theoretic treatment of the general propagator $e^{-it\mathrm{Op}^{\mathrm{w}}(a)}$ of (1), with $a(z)$ satisfying (17), in the lines of a metaplectic calculus. This allows applications to the so-called Wigner wave front set and to the boundedness in time-frequency spaces.
@@ 1.4.1. Generators
varying $E\in\operatorname*{GL}(d,\mathbb{R})$ and $Q$ real and symmetric. By allowing $E$ and $Q$ to be complex, we obtain the generators of $\operatorname*{Sp}(d,\mathbb{C})$. However, constructing a projective representation associating every $S\in\operatorname*{Sp}(d,\mathbb{C})$ to an operator bounded on $L^{2}(\mathbb{R}^{d})$, such that it agrees with the metaplectic representation on $\operatorname*{Sp}(d,\mathbb{R})$ and its image forms a group, is impossible. This also reflects the fact that, unlike the real case, problems (1) with complex quadratic Hamiltonians are generally defined only for $t\geq 0$. Indeed, $\operatorname*{Sp}_{+}(d,\mathbb{C})$ forms a semigroup rather than a group, and the question of its generators arises naturally.
@@ 1.4.2. Polar decomposition and its applications
The polar decomposition derived in [2] for the propagators of problems (1) with complex quadratic Hamiltonians is actually a very general result holding for every metaplectic operator, see Theorem 4.4. This yields the algebraic characterization of $\operatorname*{Mp}_{+}(d,\mathbb{C})$ as isomorphic to $\operatorname*{Mp}(d,\mathbb{R})\times\operatorname*{Mp}_{0}(d,\mathbb{C})$, where $\operatorname*{Mp}_{0}(d,\mathbb{C})$ is the subset of metaplectic operators that are positive and self-adjoint, and to a further characterization of the operators in $\operatorname*{Mp}_{0}(d,\mathbb{C})$ as the $\operatorname*{Mp}(d,\mathbb{R})$-conjugates of an explicit subsemigroup of metaplectic operators, here called atomic metaplectic contractions, see Section 3.2 and Corollary 4.5. Let us briefly summarize these results.
@@
can hold for $M\in\operatorname*{GL}(2d,\mathbb{R})$ if and only if the corresponding operator (possibly anti-linear) $T$ is either a metaplectic operator or the composition of a metaplectic operator with complex conjugation, as it was proven by Dias and Prata [25].
@@
Since the polar decomposition of $\widehat{S}$ is unique, formula (23) provides an unambiguous separation between the contribution of what can be regarded as a sort of real part of $\widehat{S}$, namely $\widehat{U}\in\operatorname*{Mp}(d,\mathbb{R})$, and that of its imaginary part, represented by $\widehat{Z}\in\operatorname*{Mp}_{0}(d,\mathbb{C})$. However inappropriate, these phraseologies provide a clear picture of how the positive self-adjoint operator $\widehat{Z}$, absent for metaplectic operators with real projections, affects the phase-space concentration of $\widehat{S}$. Stated differently, when $\widehat{U}\in\operatorname*{Mp}(d,\mathbb{R})$ we recover the classical symplectic covariance property of the Wigner distribution (5), whereas in the general case (23) yields an extension in which the action of $\widehat{U}$ is disentangled from the additional contribution induced by $\widehat{Z}$.